OpenStudy (anonymous):

i need help with 3n-2/5=7/10

OpenStudy (anonymous):

3n-2/5=7/10 orig question 30n-4=7 mult by 10 to eliminate the fractions. 30n=11 move to get it as one factor n=11/30 divide to get n alone

OpenStudy (anonymous):

thank you